Is there any relation between the degree of the polynomial and the number of data points in polynomial curve fitting? If yes, write the relation. If no, write a brief (no more than two lines) explanation.




1
Expert's answer
2021-11-15T19:12:41-0500

In general, whenever you have "n" data points, you can always find a polynomial of degree "n-1" which passes through all of them: 


"\begin{bmatrix}
 1 & x_1 &x_1^2 & ... & x_1^{n-1} \\
 1 & x_2 &x_2^2 & ... & x_2^{n-1} \\
 1 & x_3 &x_3^2 & ... & x_3^{n-1} \\
 \vdots & \vdots & \vdots & ... & \vdots \\
 1 & x_n &x_n^2 & ... & x_n^{n-1} \\
\end{bmatrix}\begin{bmatrix}
 c_1 \\
 c_2 \\
c_3 \\
\vdots \\
c_n 
\end{bmatrix}=\begin{bmatrix}
 y_1 \\
 y_2 \\
y_3 \\
\vdots \\
y_n 
\end{bmatrix}"